OpenStudy (casadie1004):

Help! Fan and medal. What is the ordered pair? Y=-x+2 and y=3x-1 2. What is the ordered pair? -2x+5y=-10 and -3x + 5y =-20

OpenStudy (anonymous):

for the first one, set \[-x+2=3x-1\] and solve for \(x\) in three steps

OpenStudy (anonymous):

clear what the steps are?

OpenStudy (casadie1004):

Do u add -1x from 3x or subtract?

OpenStudy (anonymous):

add \(x\) to both sides first to get rid of the \(-x\) on the left

OpenStudy (casadie1004):

I got 0.5 that's not right is it?

OpenStudy (anonymous):

no it isn't if you add \(x\) to both sides you get \[2=4x-1\] now add 1 to both sides

OpenStudy (casadie1004):

0.75, I'm sorry I forgot the 1. To find the y I just plug 0.75 in one of the first equations right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or if you like fractions, plug in \(\frac{3}{4}\) same thing

OpenStudy (casadie1004):

Lol I'm horrible at fractions. How do I do the second one?

OpenStudy (casadie1004):

Do u want me to do it on a different question so that u can get 2 medals ?

OpenStudy (anonymous):

no no this is fine here

OpenStudy (anonymous):

\[-2x+5y=-10 \\ -3x + 5y =-20\]

OpenStudy (anonymous):

multiply each term in the second equation by \(-1\)

OpenStudy (anonymous):

in other words, change the sign of everything in the second equation

OpenStudy (casadie1004):

Do u want me to do it on a different question so that u can get 2 medals ?

OpenStudy (anonymous):

no here is fine

OpenStudy (anonymous):

\[-2x+5y=-10 \\ -3x + 5y =-20\] multiply the second one by \(-1\) and you get \[-2x+5y=-10 \\ 3x - 5y =20\]

OpenStudy (anonymous):

now add, and the y terms will go bye bye

OpenStudy (casadie1004):

i didnt resend that i dont know what happened. so it becomes 1x right?

OpenStudy (anonymous):

on the left, yes, and on the right you get \(-10+20=10\)

OpenStudy (casadie1004):

what? how did u get that? I thought u would just devide 20 by one after that. How did u get the -10

OpenStudy (casadie1004):

Im sorry, im horrible at math.

OpenStudy (anonymous):

\[-2x+5y=-10 \\ 3x - 5y =20\] add in a column on the left you get just \(x\) and on the right you get \(-10+20=10\)

OpenStudy (anonymous):

which means of course that \(x=10\)

OpenStudy (anonymous):

then to find the \(y\) replace \(x\) by \(10\) in either equation

OpenStudy (casadie1004):

ohhhhhhh, never mind i see now. Im sorry, so now i just put the x in the first equation again . Thank you!

OpenStudy (anonymous):

right

OpenStudy (casadie1004):

y=2 ?

OpenStudy (anonymous):

i don't know lwet me check

OpenStudy (anonymous):

yup \(x=10,y=2\) http://www.wolframalpha.com/input/?i=-2x%2B5y%3D-10+and+-3x+%2B+5y+%3D-20

OpenStudy (casadie1004):

Thank you!

OpenStudy (anonymous):

yw

OpenStudy (casadie1004):

Hey do u know how to find out how many solutions something has?